# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v2

@@ 2.4 Workers
The value function of an employed individual with race $R$, productivity $s$, asset $a$, and working with a $np$ firm is given by $W_{np}(R,s,a)$, while that of a white worker employed with a $p$ firm is given by $W_{p}(wh,s,a)$. Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced ($U^{I}_{np}(R,s,a)$) or a prejudiced ($U^{I}_{p}(wh,s,a)$) firm previously. An unemployed worker who is not eligible for unemployment benefits has $U^{N}(R,s,a)$ over their lifetime. All the workers discount their future utility by $\beta$. Similar to Nakajima (2012) and Setty and Yedid-Levi (2021), we assume that workers cannot borrow. This imposes an exogenous constraint of $a^{\prime}\geq 0$ on all workers.
@@ 2.4.1 Employed with np firm
Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in a $np$ firm. The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their discounted lifetime utility. Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$. The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of extreme wealth shock $\epsilon_{R}$.
@@
The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3. If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$. Otherwise, one may stay employed with probability $1-\lambda_{np}$. If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
@@ 2.4.2 Employed with p firm
Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, and we only keep track of white workers since prejudiced firms do not hire black workers.
@@
The matched worker receives a job destruction shock specific to $p$ firms, with probability $\lambda_{p}$. A worker losing the job in the current period is eligible for unemployment benefits and earns value $U^{I}_{p}$ in the next period. For each $s^{\prime}$ and after realizing $\epsilon_{wh}$, an employed worker’s future value expands to Equation 5.
@@ 2.4.3 Unemployed and eligible workers
Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or a $np$ firm. In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
@@
Their future value $\hat{U^{I}}_{np}(bl,s^{\prime},a^{\prime})$ expands according to Equation 7:
@@
The unemployed worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$. Otherwise, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$. If one loses the benefit, their value function becomes $U^{N}(bl,s^{\prime},a^{\prime})$.
@@
Unemployed white workers are recruited by both $p$ and $np$ firms. Equation 8 describes their value function. Like a black worker, a white worker encounters an extreme wealth shock with probability $\epsilon_{wh}$. They also have a budget constraint from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer.
@@
Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$. With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives an only job offer from a $p$ firm. With probability $f(\theta_{np})(1-f(\theta_{p}))$, they receive an only offer from a $np$ firm. With probability $f(\theta_{np})f(\theta_{p})$, they receive offers from both $np$ and $p$ firms simultaneously. If the worker receives both $p$ and $np$ offers, they will choose the job with a higher lifetime value. The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$). In this case, one may continue to receive unemployment benefits with probability $P_{e}$. Otherwise, one moves to a state of receiving no unemployment benefit with probability $(1-P_{e})$.
@@
Equation 10 describes the problem faced by a white worker who last worked with a $p$ firm. The problem is identical to the previous case, except that they receive a payout of $b_{p}(wh,s,a)$ from the unemployment insurance.
@@
After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the unemployed worker’s future value is described in Equation 11. It is analogous to Equation 9.
@@ 2.4.4 Unemployed and ineligible workers
Ineligible workers have lost their unemployment insurance and are still searching for jobs. They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again. Therefore, their consumption relies solely on previous savings and lump-sum dividends, independent of their prior employment status.
@@
And Equation 13 expands the future value of the black worker. One has a probability of $f(\theta_{np})$ of finding a job - only with a $np$ firm; otherwise, they remain unemployed and ineligible.
@@
Equation 14 describes the problem of a white worker searching for jobs without unemployment insurance.
@@
And Equation 15 expands the future value of the white worker in this scenario. It is nearly identical to Equation 11. With a probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$, the unemployed worker does not find any job and remains unemployed and ineligible.
@@ 2.5 Firms
A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. These firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
@@ 2.5.1 Vacant np firm
Vacant $np$ firms pay a cost of $\kappa_{np}$ and search among all unemployed workers irrespective of their race. These firms match with an unemployed worker with probability $q(\theta_{np})$. Equation 16 describes the value of a vacant $np$ firm, $V_{np}$.
$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}\Big{]}$ (16)
$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}$
$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))\Big{]}\Big{\}}da,$
The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$. A $np$ firm can match with either a black or a white worker from the current unemployment pool. $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers. Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ are the densities of the unemployed black and white workers over $s$ and $a$. These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$. Since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}})$, or if the worker did not receive a $p$ firm offer, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}(1-f(\theta_{p})))$.
@@ 2.5.2 Vacant p firm
Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers. The probability of matching with an unemployed white worker is $q(\theta_{p})$. Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}$ (17)
$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))\Big{]}da$
Since the vacant $p$ firm only searches among white workers, the probability of matching with an unemployed white worker who has productivity $s$ and assets $a$ is represented by $\phi_{u}(wh,s,a)/u_{wh}$. Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than other options or if they do not receive a competing offer from an $np$ firm.
We assume that there is free entry of firms, and hence both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
@@ 2.5.3 Producing np firm
The value of a producing $np$ firm is described by Equation 18. The firm chooses capital $k$ to maximize its lifetime value producing with the worker. The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$. Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm. The future value of the firm incorporates this probability.
In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output minus production costs. The worker’s productivity, $s$, is a factor in the production function $sf(k)$. The matched firm produces $sf(k)$ units of output and incurs costs for capital rental and depreciation, represented by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
@@
2.6 Producing p firm
Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm. The firm rents capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
@@ 2.7 Wage bargaining
In each period, matched worker-firm pairs bargain for wages over the matched surplus. A worker has their matched surplus as $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ and a firm has its matched surplus as $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$. A worker’s bargaining power $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$. The resulting wage for workers at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
@@
Finally, we define the stationary equilibrium in Appendix A.
@@ 3 Calibration
Each period in the model represents a quarter. We calibrate the benchmark model to match the relevant U.S. economy moments. We have two sets of parameters. One group of parameters is chosen externally based on literature and empirical evidence without using model-generated data. The other parameters are calibrated internally by simulating our model to match a set of relevant data moments. We set the share of black workers to be 19.1% in the model, with the remaining being white workers. Table 1 shows the internally calibrated parameter values, their targeted moments, the externally chosen parameters, and their sources.
@@
$\iota$ 1.3012 matching elasticity job finding rate - black 0.4946 0.4946
$\kappa_{p}$ 4.2622 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
$\kappa_{\text{np}}$ 2.7350 np sector vacancy posting cost market tightness 1.0 1.0
$\lambda_{np}$ 0.0644 np sector job destruction shock job separation rate - black 0.0644 0.0644
$\lambda_{p}$ 0.0268 p sector job destruction shock job separation rate - white 0.0380 0.0380
$\xi_{bl}$ 0.1397 bargaining power - black mean wage ratio 0.75 0.75
$\xi_{wh}$ 0.2110 bargaining power - white firm profit share 0.033 0.033
@@
$\epsilon_{wh}$ 0.0086 extreme wealth shock - white zero wealth share - white 0.07 0.07
@@
$\chi$ 0.8433 maximum UI coverage Setty and Yedid-Levi (2021) 48% median wage
$Pe$ 0.5385 probability of UI eligibility maximum weeks of eligibility
Notes: This table reports the parameters, their values, and descriptions. The top panel presents the parameters chosen internally by minimizing the distance between model-generated moments and data. The last two columns of the top panel compare the targeted moments in the data and in the model simulations. The bottom panel reports the parameters chosen externally, their values, and description.

### v3

@@ 2.4 Workers
The value function of an employed individual with race $R$, productivity $s$, asset $a$, and working with an $np$ firm is given by $W_{np}(R,s,a)$, while that of a white worker employed with a $p$ firm is given by $W_{p}(wh,s,a)$. Since the unemployment benefit is indexed to the worker’s counterfactual wage, the values of the unemployed workers eligible for benefits depend on whether they worked with a non-prejudiced (with value function $U^{I}_{np}(R,s,a)$) or a prejudiced (with value function $U^{I}_{p}(wh,s,a)$) firm previously. An unemployed worker who is not eligible for unemployment benefits has a value of $U^{N}(R,s,a)$ over their lifetime. All the workers discount their future utility by $\beta$. Similar to Nakajima (2012) and Setty and Yedid-Levi (2021), we assume that workers cannot borrow. This imposes an exogenous constraint of $a^{\prime}\geq 0$ on all workers.
@@ 2.4.1 Employed with np firm
Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in an $np$ firm. The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their discounted lifetime utility. Their income consists of period wage, $\omega_{np}$, net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$. The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of the extreme wealth shock $\epsilon_{R}$.
@@
The worker receives a job destruction shock specific to $np$ firms and becomes unemployed with probability $\lambda_{np}$. Otherwise, they continue to remain employed with probability $1-\lambda_{np}$. If unemployed, the worker receives unemployment benefits and realizes a value of $U^{I}_{np}$. Thus, the future value of an employed worker conditional on the realization of wealth shock is given by Equation 3.
@@ 2.4.2 Employed with p firm
Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, but we only consider white workers since prejudiced firms do not hire black workers.
@@
The matched worker receives a job destruction shock specific to $p$ firms, with probability $\lambda_{p}$. A worker losing the job in the current period is eligible for unemployment benefits and earns value $U^{I}_{p}$ in the next period. Conditional on realizing the wealth shock, an employed worker’s future value expands to Equation 5.
@@ 2.4.3 Unemployed and eligible workers
Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or in an $np$ firm. In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
@@
Their future value $\hat{U}^{I}_{np}(bl,s^{\prime},a^{\prime})$ expands to Equation 7.
@@
The unemployed black worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$. Otherwise, they remain unemployed and continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$. If one loses the benefit, the worker receives a value of $U^{N}(bl,s^{\prime},a^{\prime})$.
@@
Unemployed white workers are recruited by both $p$ and $np$ firms, and their value function is given by Equation 8. Like a black worker, a white worker encounters an extreme wealth shock with probability $\epsilon_{wh}$. They also have a budget constraint depending on their post-tax unemployment insurance payout, current savings, and a lump-sum dividend transfer.
@@
Equation 9 describes the future value of an unemployed white worker conditional on the wealth shock. With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives a job offer only from a $p$ firm, while they receive an offer only from a $np$ firm with probability $f(\theta_{np})(1-f(\theta_{p}))$. With probability $f(\theta_{np})f(\theta_{p})$, they receive offers from both $np$ and $p$ firms simultaneously. If the worker receives both $p$ and $np$ offers, they will choose the job with a higher lifetime value. The white worker will remain unemployed if they don’t receive any offer (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$). In this case, the worker continues to receive unemployment benefits with probability $P_{e}$, and lose their eligibility with probability $(1-P_{e})$.
@@
Equation 10 describes the problem faced by a white worker who last worked with a $p$ firm. The problem is identical to the previous case, except that they receive an insurance payout of $b_{p}(wh,s,a)$.
@@
Analogous to the previous case, the unemployed worker’s future value conditional on the wealth shock is given by Equation 11.
@@ 2.4.4 Unemployed and ineligible workers
Ineligible workers are those who have lost their eligibility for receiving unemployment insurance and are still searching for jobs. Once ineligible, they no longer become eligible again. Therefore, their consumption relies solely on their savings and lump sum dividends, regardless of their prior employment status.
@@
Equation 13 expands the conditional future value of the black worker. The worker finds a job with an $np$ firm with probability $f(\theta_{np})$, and remains unemployed and ineligible otherwise.
@@
Similarly, Equation 14 describes the problem of a white worker searching for jobs without unemployment insurance.
@@
Equation 15 expands the conditional future value of the white worker in this scenario. It is almost identical to Equation 11 except for the fact that, in the event the unemployed worker doesn’t find any job (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$), the worker continues to remain ineligible for receiving unemployment benefits.
@@ 2.5 Firms
A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. All firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
@@ 2.5.1 Vacant np firm
Vacant $np$ firms pay a cost of $\kappa_{np}$ and search among all unemployed workers irrespective of their race. These firms match with an unemployed worker with probability $q(\theta_{np})$. The value of a vacant $np$ firm, $V_{np}$ is given by Equation 16.
$V_{np}$ $=-\kappa_{np}+\frac{q(\theta_{np})}{1+r}\int_{a}\Bigg{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{bl})J_{np}(bl,s^{\prime},g(bl,s,a))+\epsilon_{bl}J_{np}(bl,s^{\prime},0)\Big{]}\frac{\phi_{u}(bl,s,a)}{u}$ (16)
$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}(1-\epsilon_{wh})\tilde{V}_{np}(wh,s^{\prime},g(wh,s,a))+\epsilon_{wh}\tilde{V}_{np}(wh,s^{\prime},0)\Big{]}\frac{\phi_{u}(wh,s,a)}{u}\Bigg{\}}da$
where
$\tilde{V}_{np}$ $(wh,s^{\prime},g(wh,s,a))=\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))\geq W_{p}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{np}(wh,s^{\prime},g(wh,s,a))\Big{)}$
$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))<W_{p}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{np}(wh,s^{\prime},g(wh,s,a))\Big{)}(1-f(\theta_{p}))$
with $g(R,s,a)$ denoting the worker’s optimal choice of asset $a^{\prime}$ for the next period. The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the expected value of production. An $np$ firm can match with either a black or a white worker from the current unemployment pool. $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers. Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$. An unemployed black worker will accept the $np$ job offer once they get matched. On the other hand, since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, denoted by $\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))\geq W_{p}(wh,s^{\prime},g(wh,s,a))\}}$, or if the worker did not receive a $p$ firm offer, described by $\mathbbm{1}_{\{W_{np}(wh,s^{\prime},g(wh,s,a))<W_{p}(wh,s^{\prime},g(wh,s,a))\}}(1-f(\theta_{p}))$.
@@ 2.5.2 Vacant p firm
Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers. The probability of matching with an unemployed white worker is $q(\theta_{p})$. Equation 17 describes the value of maintaining a $p$ vacancy, $V_{p}$.
[…]
where
$\tilde{V}_{p}$ $(wh,s^{\prime},g(wh,s,a))=\mathbbm{1}_{\{W_{p}(wh,s^{\prime},g(wh,s,a))>W_{np}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{p}(wh,s^{\prime},g(wh,s,a))\Big{)}$
$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},g(wh,s,a))\leq W_{np}(wh,s^{\prime},g(wh,s,a))\}}\Big{(}J_{p}(wh,s^{\prime},g(wh,s,a))\Big{)}(1-f(\theta_{np}))$
Since the vacant $p$ firm searches only among white workers, the probability of matching with an unemployed white worker with productivity $s$ and asset $a$ is given by $\phi_{u}(wh,s,a)/u_{wh}$. Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than the other option or if they do not receive a competing offer from an $np$ firm.
We assume that there is free entry of firms, and therefore both $p$ and $np$ firms post vacancies until $V_{p}=0$ and $V_{np}=0$, respectively.
@@ 2.5.3 Producing np firm
The problem of a producing $np$ firm is described by Equation 18. The firm chooses capital $k$ to maximize its lifetime value, with the future value discounted using the interest rate and adjusted for the job destruction rate $\lambda_{np}$. Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm. The future value of the firm incorporates this probability.
In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output net of production costs. The matched firm produces $sf(k)$ units of output and incurs costs for renting capital and depreciation, denoted by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
@@
2.5.4 Producing p firm
Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm. The firm rents capital and discounts future profit accounting for the job destruction rate $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.
@@ 2.6 Wage bargaining
In each period, matched worker-firm pairs bargain for wages over the match surplus. A worker upon matching earns a surplus of $(W_{i}(R,s,a)-U^{I}_{i}(R,s,a))$ while the firm surplus is given by $J_{i}(Ra,s,a)$, where $i$ denotes $np$ or $p$. The bargaining power of the worker $\xi_{R}$ depends on the individual’s race, and the firm’s bargaining power is $1-\xi_{R}$. The resulting wage for workers employed at $p$ and $np$ firms are given by Equation 20 and Equation 21, respectively.
@@
We define the stationary equilibrium in Appendix A.
@@ 3 Calibration
Each period in the model represents a quarter. We calibrate the benchmark model to match the relevant US economy moments. We have two sets of parameters. One group of parameters is chosen externally based on the literature and empirical evidence without using model-generated data. The other set of parameters is calibrated internally by simulating our model to match a set of relevant data moments. Table 1 shows the values of the internal calibrated parameters, their targets, along with the externally chosen parameters and their sources. We set the percentage of black workers in the model at 14.9%, based on the Panel Study of Income Dynamics (PSID) data from 1996 to 2014.555Our sample follows Blundell et al. (2018). The share calculation follows Nakajima (2021).
@@
$\iota$ 1.3452 matching elasticity job finding rate - black 0.4946 0.4946
$\kappa_{p}$ 4.6314 p sector vacancy posting cost job finding rate - white 0.6599 0.6599
$\kappa_{\text{np}}$ 2.5534 np sector vacancy posting cost market tightness 1.0 1.0
$\lambda_{np}$ 0.0644 np sector job destruction rate job separation rate - black 0.0644 0.0644
$\lambda_{p}$ 0.0268 p sector job destruction rate job separation rate - white 0.0380 0.0380
$\xi_{bl}$ 0.1371 bargaining power - black mean wage ratio 0.75 0.75
$\xi_{wh}$ 0.1988 bargaining power - white firm profit share 0.033 0.033
@@
$\epsilon_{wh}$ 0.0088 extreme wealth shock - white zero wealth share - white 0.07 0.07
$\chi$ 0.8403 maximum UI coverage fraction of median wage 0.48 0.48
$Pe$ 0.5385 prob. of UI eligibility average weeks of eligibility 26 26
@@
Notes: This table presents the model parameters, their values, and descriptions. The top panel displays the parameters selected internally by minimizing the distance between the moments generated by the model and the moments derived from the data. The last two columns of the top panel compare the targeted moments in the data with those from the model simulations. The bottom panel provides the parameters chosen externally, their values, their descriptions and sources.